OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 18, and minor axis of length 10.

OpenStudy (anonymous):

Do you know the equation of an ellipse?

OpenStudy (anonymous):

no sir

OpenStudy (anonymous):

There is two types: Please go to this link and look at the following picture and then look at the equations below them. If you still don't get it, please tell me. http://www.mathwarehouse.com/ellipse/equation-of-ellipse.php

OpenStudy (anonymous):

ok wait yes I do and I belive its this one x squared divided by twenty five plus y squared divided by eighty one = 1

OpenStudy (anonymous):

That one will produce a minor horizontal axis of 5 and a ,major vertical axis of 9. Go to the link and look at the equation and picture. It should pop the answer to you.

OpenStudy (zale101):

\[\frac{ y^2 }{ a^2}+\frac{ x^2 }{ b^2 }=1\]

OpenStudy (zale101):

this is the form u have to use for your equation